x^3+x^2-20x=0 (this will come under cubic equation)"thank you"
Answer by nerdybill(7384)

x^3+x^2-20x=0
Start by factoring out what is common among all three terms:
x(x^2+x-20) = 0
Factor the expression inside the parenthesis:
x(x+5)(x-4) = 0
.
Set each term to zero, giving you:
x = (-5, 0, 4}